# An optimal parallel-in-time preconditioner for parabolic optimal control problems

### v1

@@ 3 Numerical Result
In this section, we provide numerical results to show the performance of our proposed preconditioners. All numerical experiments are carried out using MATLAB 2023b on a PC with Intel i5-13600KF CPU 3.50GHz and 32 GB RAM.
The CPU time in seconds is measured using MATLAB built-in functions $\mathbb{tic}$ and $\mathbb{toc}$. All Steps $1-3$ in Section 2.1 are implemented by the functions $\mathbb{dst}$ and $\mathbb{fft}$ as discrete sine transform and fast Fourier transform respectively. The GMRES solver used is implemented using the built-in functions $\mathbb{gmres}$. We choose a zero initial guess and a stopping tolerance of $10^{-6}$ based on the reduction in relative residual norms for the solver tested.
In the related tables, we denote by ‘Iter’ the number of iterations for solving a linear system by an iterative solver within the given accuracy. Denote by ‘DoF’, the number of unknowns in a linear system. Let $p^{*}$ and $y^{*}$ denote the approximate solution to $p$ and $y$, respectively. Then, we define the error measure $e_{h}$ as
$e_{h}=\left\|\begin{bmatrix}y^{*}\\ p^{*}\end{bmatrix}-\begin{bmatrix}y\\ p\end{bmatrix}\right\|_{L^{\infty}_{\tau}(L^{2}(\Omega))}.$
The time interval $[0,T]$ and the space are partitioned uniformly with the mesh step size $\tau=T/n=T/{h^{-1}}$ and $h=1/(m+1)$, respectively, where $h$ can be found in the related tables.
@@
In this example [18], we consider the following two-dimensional problem of solving (1), where $\Omega=(0,1)^{2}$, $T=1$, $a(x_{1},x_{2})=1$, and
@@
We remark that $K_{m}$ can be fast diagonalizable by the discrete sine transform matrix in this example, so we applied the fast sine transforms to solve the shifted Laplacian linear system in Step $2$ of the four-step procedures in Subsection 2.1.
@@
Table 1 presents the iteration counts, CPU time, and errors of GMRES-$\mathcal{P}_{\epsilon}$ when employing the preconditioner with various $\gamma$ values in the backward Euler method. Our observations are as follows: (i) GMRES-$\mathcal{P}_{\epsilon}$ performs excellently and consistently, maintaining stable iteration counts and CPU times across a range of $\gamma$ values; and (ii) the error decreases as the mesh is refined.
@@
Table 1: Results of GMRES-$\mathcal{P}_{\epsilon}$ for Example 3.1
@@
$\gamma$ $h$ DoF GMRES-$\mathcal{P}_{\epsilon}$
Iter CPU $e_{h}$
$10^{-10}$ $2^{-5}$ 61504 4 0.034 1.54e-2
$2^{-6}$ 508032 4 0.306 7.75e-3
$2^{-7}$ 4129024 4 2.450 3.89e-3
$2^{-8}$ 33292800 4 23.478 1.95e-3
$10^{-8}$ $2^{-5}$ 61504 6 0.044 1.54e-2
$2^{-6}$ 508032 6 0.387 7.75e-3
$2^{-7}$ 4129024 6 3.589 3.89e-3
$2^{-8}$ 33292800 7 32.822 1.95e-3
$10^{-6}$ $2^{-5}$ 61504 8 0.054 1.54e-2
$2^{-6}$ 508032 10 0.740 7.71e-3
$2^{-7}$ 4129024 10 6.126 3.86e-3
$2^{-8}$ 33292800 12 60.325 1.93e-3
$10^{-4}$ $2^{-5}$ 61504 11 0.074 1.42e-2
$2^{-6}$ 508032 11 0.739 7.09e-3
$2^{-7}$ 4129024 9 5.124 3.56e-3
$2^{-8}$ 33292800 6 31.904 1.78e-3
$10^{-2}$ $2^{-5}$ 61504 12 0.075 3.10e-3
$2^{-6}$ 508032 12 0.882 1.50e-3
$2^{-7}$ 4129024 14 8.125 7.40e-4
$2^{-8}$ 33292800 14 68.417 3.67e-4
$1$ $2^{-5}$ 61504 8 0.066 7.19e-4
$2^{-6}$ 508032 8 0.531 3.65e-4
$2^{-7}$ 4129024 8 4.665 1.84e-4
$2^{-8}$ 33292800 8 38.146 9.25e-5
@@
In this example, we consider the following two-dimensional problem of solving (1) with a variable function $a(x_{1},x_{2})$, where $\Omega=(0,1)^{2}$, $T=1$, $a(x_{1},x_{2})=10^{-5}\sin{(\pi x_{1}x_{2})}$, and
@@
Since $K_{m}$ is not fast diagonalizable in this example, we applied one iteration of the V-cycle geometric multigrid method to solve the shifted Laplacian linear system (as detailed in Subsection 2.1). The Gauss-Seidel method is employed as the pre-smoother for the multigrid method.
@@
Table 2 presents the iteration counts, CPU time, and errors for GMRES-$\mathcal{P}_{\epsilon}$ when the preconditioner is utilized with various values of $\gamma$ in the backward Euler method. The purpose of this example is to evaluate the effectiveness of our solvers when the coefficient function $a(x_{1},x_{2})$ is non-constant.
@@
Similar to the last example, the results indicate that (i) GMRES-$\mathcal{P}_{\epsilon}$ achieves stable iteration counts and CPU times across a broad range of $\gamma$ values; (ii) the errors decrease as expected as the mesh refines.
@@
Table 2: Results of GMRES-$\mathcal{P}_{\epsilon}$ for Example 3.2

### v2

@@ 3 Numerical Result
In this section, we provide numerical results to show the performance of our proposed preconditioners. All numerical experiments are carried out using MATLAB R2024b on a MacBook Pro equipped with Apple M4 Pro and 24GB RAM..
The CPU time in seconds is measured using MATLAB built-in functions $\mathbb{tic}$ and $\mathbb{toc}$. All Steps $1-3$ in Section 2.1 are implemented by the functions $\mathbb{dst}$ and $\mathbb{fft}$ as discrete sine transform and fast Fourier transform respectively. The GMRES solver used is implemented using the built-in function $\mathbb{gmres}$. We choose a zero initial guess and a stopping tolerance of $10^{-8}$, based on the reduction in relative residual norms observed for the solver tested.
In the related tables, we denote by ‘Iter’ the number of iterations for solving a linear system by an iterative solver within the given accuracy. Denote with ‘DoF’, the number of unknowns in a linear system. Also, the notation GMRES-$\mathcal{P}_{\epsilon}/\mathcal{P}$ denotes the GMRES solver applied with our proposed preconditioner $\mathcal{P}{{}_{\epsilon}}$ or the original RBD preconditioner $\mathcal{P}$ proposed in [4].
The time interval $[0,T]$ and the spatial domain are partitioned uniformly with the mesh step size $\tau=T/(n+1)$ and $h=1/(m+1)$, respectively, where $h$ can be found in the corresponding tables. Note that a finite difference method is used throughout the numerical tests.
We define the error measure $e_{h}$ as
$e_{h}$ $:=\left\|\begin{bmatrix}\mathbf{y}\\ \mathbf{p}\end{bmatrix}-\begin{bmatrix}y\\ p\end{bmatrix}\right\|_{L^{\infty}_{\tau}(L^{2}(\Omega))}$
$:=h^{d/2}\max\limits_{1\leq j\leq N}\max\left\{||{\bf y}_{m}^{(j)}-y(\Omega_{h},j)||_{2},||{\bf p}_{m}^{(j-1)}-p(\Omega_{h},j-1)||_{2}\right\},$
where ${\bf y}_{m}^{(j)}$ (${\bf p}_{m}^{(j-1)}$, respectively) denotes the numerical solution of $y$ ($p$, respectively) at $j$-th ($j-1$-th, respectively) time step; $y(\Omega_{h},j)$ ($p(\Omega_{h},j-1)$, respectively) denote the values of analytical solution of $y$ ($p$, respectively) on the spatial grid at $j$-th ($j-1$-th, respectively) time step.
@@
In this example [21], we consider the following two-dimensional problem , which involves solving (1) with $\Omega=(0,1)^{2}$, $T=1$, $a(x_{1},x_{2})=1$, and
@@
We remark that $K_{m}$ can be fast diagonalizable by the discrete sine transform matrix in this example, so we applied the fast sine transforms to solve the shifted Laplacian linear systems in Step $2$ of the four-step procedures in Subsection 2.1.
@@
Tables 1 & 2 present the iteration counts, CPU times, and discretization errors of GMRES-$\mathcal{P}_{\epsilon}$ when employing the preconditioner with various $\gamma$ values in the backward Euler method. Our observations are as follows: (i) GMRES-$\mathcal{P}_{\epsilon}$ performs excellently and consistently, maintaining stable iteration counts and CPU times across a range of $\gamma$ values; and (ii) the error decreases as the mesh is refined.
To further illustrate the effectiveness of our proposed solver compared to an existing one, we present the numerical results of GMRES-$\mathcal{P}{\epsilon}$ and GMRES-$\mathcal{P}$ in Tables 3 and 4. Throughout all tests in these tables, the mesh size is fixed at $h=2^{-5}$. In virtually all cases, as $\tau$ decreases, the corresponding iteration counts for both methods stabilize. Moreover, the CPU time required by our proposed GMRES-$\mathcal{P}{\epsilon}$ is significantly lower when $\tau$ decreases, demonstrating its superior efficiency compared to GMRES-$\mathcal{P}$.
Figure 1 presents the plot of the residual ratio, defined as $\frac{\|\mathbf{r}_{j}\|_{2}}{\|\mathbf{r}_{j-1}\|_{2}}$, against the iteration number $j$ required for convergence under various regularization parameters. The results demonstrate that the theoretically optimal upper bound derived in Theorem 2.21, given by
$\lim_{\delta\to 0^{+}}\frac{\sqrt{-\delta^{2}+8\delta+2}}{2+\delta}=\frac{\sqrt{2}}{2},$
can indeed be achieved. Moreover, in some instances, the residual decreases even more rapidly than this theoretical estimate suggests.
@@
Table 1: Results of GMRES-$\mathcal{P}_{\epsilon}$ for Example 3.1 with $\gamma=10^{-10},10^{-8},10^{-6}$.
$\gamma$ $h$ $\tau$ DoF GMRES-$\mathcal{P}_{\epsilon}$
Iter CPU $e_{h,\tau}$
$10^{-10}$ $2^{-6}$ $2^{-6}$ 508032 6 0.23 2.60e-05
$2^{-7}$ 1016064 6 0.40 9.81e-06
$2^{-8}$ 2032128 6 0.72 3.95e-06
$2^{-7}$ $2^{-6}$ 2064512 6 0.81 1.94e-05
$2^{-7}$ 4129024 6 1.65 6.11e-06
$2^{-8}$ 8258048 6 3.48 1.97e-06
$2^{-8}$ $2^{-6}$ 8323200 6 3.35 1.78e-05
$2^{-7}$ 16646400 6 7.26 5.18e-06
$2^{-8}$ 33292800 6 16.17 1.47e-06
$10^{-8}$ $2^{-6}$ $2^{-6}$ 508032 8 0.28 2.60e-05
$2^{-7}$ 1016064 8 0.53 9.81e-06
$2^{-8}$ 2032128 10 1.28 3.95e-06
$2^{-7}$ $2^{-6}$ 2064512 8 1.04 1.94e-05
$2^{-7}$ 4129024 8 2.07 6.11e-06
$2^{-8}$ 8258048 10 5.36 1.97e-06
$2^{-8}$ $2^{-6}$ 8323200 8 4.10 1.78e-05
$2^{-7}$ 16646400 8 8.87 5.19e-06
$2^{-8}$ 33292800 10 25.98 1.47e-06
$10^{-6}$ $2^{-6}$ $2^{-6}$ 508032 12 0.49 2.61e-05
$2^{-7}$ 1016064 14 1.06 9.99e-06
$2^{-8}$ 2032128 17 2.45 4.21e-06
$2^{-7}$ $2^{-6}$ 2064512 12 1.54 1.95e-05
$2^{-7}$ 4129024 14 3.71 6.22e-06
$2^{-8}$ 8258048 17 9.66 2.10e-06
$2^{-8}$ $2^{-6}$ 8323200 12 6.17 1.79e-05
$2^{-7}$ 16646400 14 15.49 5.28e-06
$2^{-8}$ 33292800 17 51.43 1.57e-06
@@
Table 2: Results of GMRES-$\mathcal{P}_{\epsilon}$ for Example 3.1 with $\gamma=10^{-4},10^{-2},1$.
$\gamma$ $h$ $\tau$ DoF GMRES-$\mathcal{P}_{\epsilon}$
Iter CPU $e_{h,\tau}$
$10^{-4}$ $2^{-6}$ $2^{-6}$ 508032 15 0.65 3.59e-05
$2^{-7}$ 1016064 13 0.96 1.92e-05
$2^{-8}$ 2032128 7 0.95 1.26e-05
$2^{-7}$ $2^{-6}$ 2064512 15 2.03 2.69e-05
$2^{-7}$ 4129024 13 3.48 1.19e-05
$2^{-8}$ 8258048 7 3.93 6.25e-06
$2^{-8}$ $2^{-6}$ 8323200 15 8.01 2.46e-05
$2^{-7}$ 16646400 13 14.46 1.01e-05
$2^{-8}$ 33292800 7 18.77 4.68e-06
$10^{-2}$ $2^{-6}$ $2^{-6}$ 508032 17 0.85 2.02e-04
$2^{-7}$ 1016064 18 1.48 1.36e-03
$2^{-8}$ 2032128 18 2.63 1.03e-03
$2^{-7}$ $2^{-6}$ 2064512 17 2.41 1.51e-04
$2^{-7}$ 4129024 18 4.78 8.49e-05
$2^{-8}$ 8258048 18 9.92 5.13e-05
$2^{-8}$ $2^{-6}$ 8323200 17 8.94 1.38e-04
$2^{-7}$ 16646400 18 19.84 7.20e-05
$2^{-8}$ 33292800 18 52.71 3.83e-05
$1$ $2^{-6}$ $2^{-6}$ 508032 10 0.38 2.49e-03
$2^{-7}$ 1016064 10 0.69 1.68e-03
$2^{-8}$ 2032128 10 1.28 1.27e-04
$2^{-7}$ $2^{-6}$ 2064512 10 1.25 1.86e-04
$2^{-7}$ 4129024 10 2.62 1.04e-04
$2^{-8}$ 8258048 10 5.44 6.31e-05
$2^{-8}$ $2^{-6}$ 8323200 10 5.28 1.70e-04
$2^{-7}$ 16646400 10 11.21 8.86e-05
$2^{-8}$ 33292800 10 25.21 4.72e-05
Table 3: Results of GMRES-$\mathcal{P}_{\epsilon}$ for Example 3.1 with various regularization parameters $\gamma$
$\gamma$ $h$ $\tau$ DoF GMRES-$\mathcal{P}_{\epsilon}$
Iter CPU $e_{h,\tau}$
$10^{-10}$ $2^{-5}$ $2^{-5}$ 61504 6 0.05 1.13e-04
$2^{-6}$ 123008 6 0.08 5.21e-05
$2^{-7}$ 246016 6 0.15 2.46e-05
$2^{-8}$ 492032 6 0.24 1.19e-05
$2^{-9}$ 984064 6 0.44 5.82e-06
$2^{-10}$ 1968128 8 0.98 2.88e-06
$2^{-11}$ 3936256 8 2.12 1.43e-06
$10^{-8}$ $2^{-5}$ $2^{-5}$ 61504 6 0.06 1.13e-04
$2^{-6}$ 123008 8 0.16 5.21e-05
$2^{-7}$ 246016 8 0.23 2.46e-05
$2^{-8}$ 492032 10 0.47 1.19e-05
$2^{-9}$ 984064 11 0.84 5.84e-06
$2^{-10}$ 1968128 13 1.78 2.91e-06
$2^{-11}$ 3936256 16 4.20 1.49e-06
$10^{-6}$ $2^{-5}$ $2^{-5}$ 61504 10 0.09 1.13e-04
$2^{-6}$ 123008 12 0.22 5.24e-05
$2^{-7}$ 246016 14 0.43 2.51e-05
$2^{-8}$ 492032 17 0.89 1.27e-05
$2^{-9}$ 984064 18 1.43 7.12e-06
$2^{-10}$ 1968128 17 2.27 4.75e-06
$2^{-11}$ 3936256 15 3.92 3.75e-06
$10^{-4}$ $2^{-5}$ $2^{-5}$ 61504 15 0.16 1.29e-04
$2^{-6}$ 123008 15 0.32 7.20e-05
$2^{-7}$ 246016 13 0.42 4.81e-05
$2^{-8}$ 492032 7 0.30 3.78e-05
$2^{-9}$ 984064 12 0.95 3.32e-05
$2^{-10}$ 1968128 13 1.79 3.10e-05
$2^{-11}$ 3936256 15 3.92 2.99e-05
$10^{-2}$ $2^{-5}$ $2^{-5}$ 61504 16 0.20 5.30e-04
$2^{-6}$ 123008 17 0.33 4.06e-04
$2^{-7}$ 246016 18 0.64 3.42e-04
$2^{-8}$ 492032 18 0.95 3.10e-04
$2^{-9}$ 984064 18 1.43 2.94e-04
$2^{-10}$ 1968128 19 2.75 2.86e-04
$2^{-11}$ 3936256 19 5.11 2.82e-04
$1$ $2^{-5}$ $2^{-5}$ 61504 10 0.12 6.52e-04
$2^{-6}$ 123008 10 0.18 4.99e-04
$2^{-7}$ 246016 10 0.27 4.21e-04
$2^{-8}$ 492032 10 0.46 3.82e-04
$2^{-9}$ 984064 10 0.77 3.62e-04
$2^{-10}$ 1968128 10 1.47 3.52e-04
$2^{-11}$ 3936256 10 2.96 3.47e-04
Table 4: Results of GMRES-$\mathcal{P}$ for Example 3.1 with various regularization parameters $\gamma$
$\gamma$ $h$ $\tau$ DoF GMRES-$\mathcal{P}$
Iter CPU $e_{h,\tau}$
$10^{-10}$ $2^{-5}$ $2^{-5}$ 61504 6 0.14 1.13e-04
$2^{-6}$ 123008 6 0.19 5.21e-05
$2^{-7}$ 246016 6 0.37 2.46e-05
$2^{-8}$ 492032 6 0.70 1.19e-05
$2^{-9}$ 984064 6 1.36 5.82e-06
$2^{-10}$ 1968128 8 3.59 2.88e-06
$2^{-11}$ 3936256 8 7.27 1.43e-06
$10^{-8}$ $2^{-5}$ $2^{-5}$ 61504 6 0.13 1.13e-04
$2^{-6}$ 123008 8 0.28 5.21e-05
$2^{-7}$ 246016 8 0.52 2.46e-05
$2^{-8}$ 492032 10 1.18 1.19e-05
$2^{-9}$ 984064 11 2.53 5.84e-06
$2^{-10}$ 1968128 13 6.08 2.91e-06
$2^{-11}$ 3936256 16 13.53 1.49e-06
$10^{-6}$ $2^{-5}$ $2^{-5}$ 61504 10 0.17 1.13e-04
$2^{-6}$ 123008 12 0.41 5.24e-05
$2^{-7}$ 246016 14 0.92 2.51e-05
$2^{-8}$ 492032 17 2.28 1.27e-05
$2^{-9}$ 984064 18 4.45 7.12e-06
$2^{-10}$ 1968128 17 7.41 4.75e-06
$2^{-11}$ 3936256 15 12.96 3.75e-06
$10^{-4}$ $2^{-5}$ $2^{-5}$ 61504 15 0.29 1.29e-04
$2^{-6}$ 123008 15 0.54 7.20e-05
$2^{-7}$ 246016 13 0.91 4.81e-05
$2^{-8}$ 492032 7 0.90 3.78e-05
$2^{-9}$ 984064 12 3.06 3.32e-05
$2^{-10}$ 1968128 13 6.16 3.10e-05
$2^{-11}$ 3936256 15 13.52 2.99e-05
$10^{-2}$ $2^{-5}$ $2^{-5}$ 61504 16 0.31 5.30e-04
$2^{-6}$ 123008 17 0.66 4.06e-04
$2^{-7}$ 246016 18 1.28 3.42e-04
$2^{-8}$ 492032 18 2.53 3.10e-04
$2^{-9}$ 984064 18 4.70 2.94e-04
$2^{-10}$ 1968128 19 9.20 2.86e-04
$2^{-11}$ 3936256 19 17.38 2.82e-04
$1$ $2^{-5}$ $2^{-5}$ 61504 10 0.18 6.52e-04
$2^{-6}$ 123008 10 0.38 4.99e-04
$2^{-7}$ 246016 10 0.66 4.21e-04
$2^{-8}$ 492032 10 1.25 3.82e-04
$2^{-9}$ 984064 10 2.44 3.62e-04
$2^{-10}$ 1968128 10 5.07 3.52e-04
$2^{-11}$ 3936256 10 10.06 3.47e-04
Figure 1: Convergence factor per iteration for Example 3.1 with various regularization parameters $\gamma$.
@@
In this example, we consider the following two-dimensional problem , which involves solving (1) with a variable function $a(x_{1},x_{2})$, $\Omega=(0,1)^{2}$, $T=1$, $a(x_{1},x_{2})=10^{-5}\sin{(\pi x_{1}x_{2})}$, and
@@
Since $K_{m}$ is not fast diagonalizable in this example, we applied one iteration of the V-cycle geometric multigrid method to solve the shifted Laplacian linear system (as detailed in Subsection 2.1). The Gauss-Seidel method is employed as the pre-smoother for the multigrid method , and no post-smoother is used.
@@
Tables 5 & 6 present the iteration counts, CPU times, and discretization errors for GMRES-$\mathcal{P}_{\epsilon}$ when the preconditioner is utilized with various values of $\gamma$ in the backward Euler method. The purpose of this example is to evaluate the effectiveness of our solvers when the coefficient function $a(x_{1},x_{2})$ is non-constant.
Similar to the previous example, we present the numerical results of GMRES-$\mathcal{P}{\epsilon}$ and GMRES-$\mathcal{P}$ in Tables 7 and 8 for comparison purposes. Throughout all tests reported in these tables, the mesh size is fixed at $h=2^{-5}$. In nearly all cases, as the time-step size $\tau$ decreases, the iteration counts for both methods tend to stabilize. Furthermore, the CPU time required by our proposed GMRES-$\mathcal{P}_{\epsilon}$ solver is consistently lower for smaller values of $\tau$, showcasing better computational performance than GMRES-$\mathcal{P}$.
@@
Similar to the last example, the results indicate that (i) GMRES-$\mathcal{P}_{\epsilon}$ achieves stable iteration counts and CPU times across a broad range of $\gamma$ values; (ii) the error decreases as expected as the mesh refines.
Similarly to Figure 1, Figure 2 illustrates the residual ratio $\frac{\|\mathbf{r}_{j}\|_{2}}{\|\mathbf{r}_{j-1}\|_{2}}$ plotted against the iteration number $j$ for various regularization parameters. The convergence behavior remains consistent with the theoretical upper bound from Theorem 2.21, where the optimal upper bound $\frac{\sqrt{2}}{2}$ is attained. In fact, as observed previously, the residual reduction occasionally outpaces this theoretical estimate.
@@
Table 5: Results of GMRES-$\mathcal{P}_{\epsilon}$ for Example 3.2 with $\gamma=10^{-10},10^{-8},10^{-6}$
$\gamma$ $h$ $\tau$ DoF GMRES-$\mathcal{P}_{\epsilon}$
Iter CPU $e_{h,\tau}$
$10^{-10}$ $2^{-6}$ $2^{-6}$ 508032 6 0.35 1.50e-06
$2^{-7}$ 1016064 6 0.66 3.75e-07
$2^{-8}$ 2032128 6 1.30 9.37e-08
$2^{-7}$ $2^{-6}$ 2064512 6 1.43 1.50e-06
$2^{-7}$ 4129024 6 2.58 3.75e-07
$2^{-8}$ 8258048 6 5.01 9.37e-08
$2^{-8}$ $2^{-6}$ 8323200 6 5.14 1.50e-06
$2^{-7}$ 16646400 6 10.36 3.75e-07
$2^{-8}$ 33292800 6 21.24 9.37e-08
$10^{-8}$ $2^{-6}$ $2^{-6}$ 508032 8 0.39 1.50e-06
$2^{-7}$ 1016064 8 0.77 3.75e-07
$2^{-8}$ 2032128 10 1.95 9.37e-08
$2^{-7}$ $2^{-6}$ 2064512 8 1.70 1.50e-06
$2^{-7}$ 4129024 8 3.25 3.75e-07
$2^{-8}$ 8258048 10 7.76 9.37e-08
$2^{-8}$ $2^{-6}$ 8323200 8 6.38 1.50e-06
$2^{-7}$ 16646400 8 12.87 3.75e-07
$2^{-8}$ 33292800 10 33.30 9.37e-08
$10^{-6}$ $2^{-6}$ $2^{-6}$ 508032 12 0.62 1.51e-06
$2^{-7}$ 1016064 15 1.52 3.81e-07
$2^{-8}$ 2032128 18 3.69 9.94e-08
$2^{-7}$ $2^{-6}$ 2064512 12 2.54 1.51e-06
$2^{-7}$ 4129024 15 5.94 3.81e-07
$2^{-8}$ 8258048 18 13.32 9.94e-08
$2^{-8}$ $2^{-6}$ 8323200 12 9.27 1.51e-06
$2^{-7}$ 16646400 15 23.49 3.81e-07
$2^{-8}$ 33292800 18 58.45 9.94e-08